OpenStudy (anonymous):

integral of 1/(cos (x) -1)

OpenStudy (anonymous):

can we rewrite this as secx-1

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{\cos(x)-1} \cdot \frac{\cos(x)+1}{\cos(x)+1} dx=\int\limits_{}^{}\frac{\cos(x)+1}{\cos^2(x)-1} dx\] \[\int\limits_{}^{}\frac{\cos(x)+1}{-(1-\cos^2(x))}dx=\int\limits_{}^{}\frac{\cos(x)+1}{-\sin^2(x)} dx\] \[\int\limits_{}^{}\frac{\cos(x)}{-\sin^2(x)} dx +\int\limits_{}^{}\frac{1}{-\sin^2(x)} dx\]

myininaya (myininaya):

\[-\int\limits_{}^{}\frac{\cos(x)}{\sin^2(x)} dx-\int\limits_{}^{}\frac{1}{\frac{1}{2}(1-\cos(2x)} dx\]

myininaya (myininaya):

for the first one let u=sin(x) let me know if you need more help

OpenStudy (anonymous):

oh, nice job myin, multiplying by the conjugate

OpenStudy (anonymous):

how do you know this???

OpenStudy (anonymous):

so what is the actual answer?

OpenStudy (anonymous):

and just to show off a little if you want \[\int\frac{1}{-\sin^2(x)}dx\] i would just recall that this is \[-\int\csc^2(x)dx\] so that the integrand is instantly the derivative of \[\cot(x)\]

OpenStudy (anonymous):

i get an actual answer of \[\csc(x)+\cot(x)\]

OpenStudy (anonymous):

that is the actual answer

OpenStudy (anonymous):

in actuality

myininaya (myininaya):

how do i know this?

OpenStudy (anonymous):

that was my question. it was a serious one. i would never know to do this

myininaya (myininaya):

well i wanted to get the bottom as one term so i can break it up

OpenStudy (anonymous):

although i have to say i like my method of solving \[-\int\frac{1}{\sin^2(x)}dx\]

myininaya (myininaya):

yes i wasn't thinking there

OpenStudy (anonymous):

ah that makes sense.

OpenStudy (anonymous):

also might explain why we see all those trig problems that say "show \[\frac{\sin(x)}{1+\cos(x)}=\text{whatever}\]

myininaya (myininaya):

oh yes identities

myininaya (myininaya):

it can be useful in trig

myininaya (myininaya):

i mean in calc

OpenStudy (anonymous):

guess they prove useful later. only problem is it is two semesters later so they forget

myininaya (myininaya):

right

myininaya (myininaya):

some people don't even take trig

OpenStudy (anonymous):

same as those problems that say find \[\tan(\sin^{-1}(x))\] by the time they see it again it is in calc 2

myininaya (myininaya):

i love those problems

OpenStudy (anonymous):

i know you do. actually i do too

myininaya (myininaya):

i was like telling my trig students make sure you know how to do these i love these

OpenStudy (anonymous):

try to explain as best i can that is it exactly the same as those problems that say "find the other trig functions of theta if sin(theta) = whatever

myininaya (myininaya):

thats key for its gonna be on the test

OpenStudy (anonymous):

ti 89 does it symbolically.

OpenStudy (anonymous):

lol ok gnight

myininaya (myininaya):

goodnight